# A Fast Parallel Approach for Neighborhood-based Link Prediction by Disregarding Large Hubs

### v1

@@ Abstract.
Link prediction can help rectify inaccuracies in community detection stemming from unaccounted-for or overlooked links within networks. Many existing works use a baseline approach, which incurs unnecessary computational costs due to its high time complexity. Further, many studies focus on smaller graphs, which can lead to misleading conclusions. The report introduces two parallel approaches, called IHub and LHub, which predict links using neighborhood-based similarity measures on large graphs. LHub is a heuristic approach, which additionally disregards large hubs — based on the idea that low-degree nodes contribute significant similarity among neighbors. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, LHub is on average 563×563$\times$563 × faster than IHub, especially on web graphs and social networks, while having similar prediction accuracy. Notably, LHub achieves a link prediction rate of $38.1M$ edges/s and improves performance at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1.1. Our Contributions
In this report, we introduced two novel approaches, namely IHub and LHub, for efficient link prediction in large graphs. The IHub approach is an enhanced parallel method that efficiently finds common neighbors and handles large graphs by tracking top-$k$ edges per-thread and later merging them globally. LHub is a heuristic approach which additionally discards large hubs, based on the observation that low-degree nodes contribute significant similarity among neighbors. We then experimentally determined suitable hub limits, i.e, the degree above which a vertex is considered as a large hub, for link prediction with each similarity metric.
On a machine with two 16-core Intel Xeon Gold 6226R processors, our results show that the LHub approach outperforms IHub by over 963×963$\times$963 × and 163×163$\times$163 × on average with $10^{-2}|E|$ and $0.1|E|$ unobserved edges, respectively. This speedup is achieved while maintaining comparable F1 scores. Notably, LHub achieves a link prediction rate of $38.1M$ edges/s with $0.1|E|$ unobserved edges. We also identified suitable similarity metric for each type of graph. When predicting $0.1|E|$ edges with the LHub approach, we observed that $63\%$ of the runtime is spent on the scoring phase, and especially higher on social networks, with high average degree. With doubling of threads, LHub exhibits an average performance scaling of 1.6×1.6$\times$1.6 ×.
@@ 4. Approach
4.1. Our Improved Approach (IHub)
@@ 4.1.3. Design of Hashtable
(a) Standard approach (IHub)
(b) Disregard hubs with degree $>8$ (LHub)
(c) Disregard hubs with degree $>4$ (LHub)
Figure 1. Illustration of our link prediction approach which disregards large hubs, i.e., first-order neighbors with high degree. Here we focus on the neighborhood of vertex $1$, but the approach applies to each vertex in the graph. In the figure, $1$ is outlined in black, its first-order neighbors in red, its second-order neighbors in blue, and explored/traversed vertices are shown with a yellow fill. Edge directions indicate traversal, with some second order vertices omitted for simplicity (dotted edges). (a) Depicts the standard approach (IHub), which considers all second-order neighbors of vertex $1$. (b) Presents our LHub approach, which considers only second-order neighbors linked to $1$ through a small hub (degree $\leq 8$). This pruning reduces runtime and enhances prediction quality. (c) Illustrates LHub approach, where vertices with degree $>4$ are considered large hubs.
@@ 4.1.7. Parallelizing the process
We call this approach IHub. It has a time complexity of $O(ND^{2})$, where $N$ is the number of vertices in the graph, and $D$ is the maximum degree of a vertex in the graph.
4.2. Our Heuristic Approach which Disregards Large Hubs (LHub)
@@ 4.2.1. Explanation of the approach
We call this approach LHub. It has the same time complexity as IHub, i.e., $O(ND^{2})$, but significantly outperforms it by runtime, while achieving the same or better prediction quality.
@@ 4.2.2. A simple example
Figure 1 demonstrates the LHub approach. Here, we consider the neighborhood of vertex $1$ in a graph, where $1$ is outlined in black, its first-order neighbors in red, its second-order neighbors in blue, and explored/traversed vertices are shown with a yellow fill. The arrows in the figure indicate the direction of DFS traversal, and for figure simplicity, three neighbors of vertex $2$ are omitted (shown with dotted edges).
Figure 1(a) depicts the standard approach, which considers all second-order neighbors of vertex $1$ for score computation, i.e. vertices $6$ to $12$. This is the process followed by the IHub approach.
Figure 1(b) depicts our LHub approach. It considers only second-order neighbors linked to vertex $1$ by a small hub, with a degree $\leq 8$, for score computation — i.e., vertices $9$ to $12$ which are linked to $1$ through vertices $3$, $4$, and $5$. It is based on the idea, as mentioned above, that low-degree vertices contribute more similarity among their neighbors, while vertices with high degree do not. Note that second-order neighbors of $1$ can still be considered if linked to $1$ by a small hub (e.g., vertex $9$ linked to $1$ through $3$).
Lastly, Figure 1(c) shows the LHub approach, considering only second-order neighbors linked to vertex $1$ by a small hub with a degree $\leq 4$ ($11$ and $12$ linked to $1$ through $4$ and $5$). First order neighbors of $1$ with a degree $>4$ are considered large hubs, and thus their second-order neighbors are disregarded.
@@ 4.2.3. Non-determinism in the result
It’s worth noting that the results obtained by our parallel algorithms (IHub and LHub) are non-deterministic. This is because multiple edges can have matching scores, and the order of edges with the same score may be random due to parallelism.
@@ 4.3. Selecting Suitable Hub Limit for each Similarity Metric
We now select a suitable hub limit $L_{H}$, i.e., a degree above which a vertex is considered a large hub, for each similarity metric presented in Section 3.1. For this, we adjust the hub limit $L_{H}$ from $2$ to $1024$ (in multiples of $2$), for each similarity metric, and adjust the number of unobserved edges $E^{U}$ from $10^{-4}|E|$ to $0.1|E|$ on a number of graphs. We then apply each similarity metric based link prediction to predict $N_{P}=|E^{U}|$ edges with the highest similarity scores. We also test with a hub limit $L_{H}$ of $\infty$, which is the IHub approach.
Figure 3(a) shows the mean runtime taken for each similarity metric, with each hub limit $L_{H}$, and with the number of unobserved edges $E^{U}$ ranging from $10^{-4}|E|$ to $0.1|E|$, while Figure 3(b) shows the mean F1 score of the predicted edges, with the IHub and LHub approaches. Results indicate that selecting a lower hub limit $L_{H}$ decreases the overall runtime of the LHub approach. In terms of both F1 score and relative runtime, we observe that a hub limit $L_{H}$ of $4$ is suitable for HP and LHN metrics, a hub limit $L_{H}$ of $32$ is suitable for CN and AA metrics, and a hub limit $L_{H}$ of $256$ is suitable for the remaining metrics (i.e., JC, SI, SC, HD, and RA). These hub limits are highlighted with thick lines in the figures. Further, as Figure 7 shows, hub limits $L_{H}$ of $4$, $32$, and $256$ offer a mean speedup of 67×67$\times$67 ×, 32×32$\times$32 ×, and 13×13$\times$13 × when the number of unobserved edges $E^{U}$ is $0.1|E|$. Indeed, disregarding large hubs, with the LHub approach, offers a significant improvement in performance with little to no degradation in the accuracy of link prediction.
4.4. Our implementation of Neighborhood-based Link Prediction
We now explain the implementation of our LHub approach for parallel neighborhood-based link prediction, which disregards large hubs. The pseudocode for this is given in Algorithm 1. Here, the predictLinks function accepts the input graph $G$, the maximum number of edges to predict $N_{P}$, a threshold score $S_{th}$, and outputs the list of predicted edges $P$. The algorithm operates in two phases: the scoring phase and the merging phase.
@@ 4.4. Our implementation of Neighborhood-based Link Prediction
(a) Relative runtime (logarithmic scale), of each link prediction method
(b) F1 score of predicted links (logarithmic scale), of each link prediction method
Figure 3. Impact of adjusting the hub limit $L_{H}$ from $2$ to $1024$ (in multiples of $2$), and to $\infty$, on the runtime (in seconds, log-scale), and precision of predicted links (in percentage, log scale), of each neighbor-based link prediction method, with the number of unobserved edges $E^{U}$ ranging from $10^{-4}|E|$ to $0.1|E|$. The full form of each link prediction method is given in Section 3.1.
Algorithm 1 Our Parallel Link Prediction algorithm.
@@ 5.1.3. Dataset
Table 1. List of $13$ graphs obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019), with the directed graphs being marked with an asterisk. Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges and removing self-loops), and $D_{avg}$ is the average degree. TOFIX $|E|$.
@@ 5.2. Comparative Performance Evaluation
In this section, we compare the performance of the LHub approach, which disregards large hubs, with the IHub approach (which does not). We conduct this comparison for observed graphs $E^{O}$ based on each graph in the dataset, with $10^{-2}|E|$ to $0.1|E|$ unobserved edges. This involves removing $10^{-2}|E|$ to $0.1|E|$ edges, with endpoints chosen uniformly at random (as explained in Section 5.1.4). We predict the same number of edges with both approaches, i.e., $N_{P}=10^{-2}|E|$ or $0.1|E|$. For each observed graph, we predict links using the IHub approach and the LHub approach with a suitable hub limit $L_{H}$ identified in Section 4.3. In Figures 4(a), 4(b), and 4(c), we plot the runtimes, speedups, and F1 scores, respectively, for only the best approach for each graph (considering both F1 score and runtime). In the figures, the labels indicate the abbreviations of the similarity metric used, followed by the value of the hub limit $L_{H}$ parameter setting (e.g., the label $CN32$ stands for the Common Neighbors (CN) similarity metric, with a hub limit $L_{H}$ of $32$, where first order neighbors with a degree greater than $32$ are avoided). It’s worth noting that a hub limit $L_{H}$ of $\infty$ essentially represents the IHub approach. Note that we were unable to run the IHub approach on large web graphs, such as webbase-2001, it-2004, and sk-2005, due to high runtimes, and their plots are not shown.
@@
(a) Runtime in seconds (logarithmic scale) for link prediction using the best similarity measure, with IHub and LHub approaches
@@
(b) Speedup (logarithmic scale) for link prediction with the best similarity measure of LHub approach, compared to IHub approach
@@
(c) F1 score of predicted links (logarithmic scale), for link prediction using the best similarity measure, with IHub and LHub approaches
@@
Figure 4. Runtime in seconds (log-scale), speedup (log-scale), and F1 score of predicted links (log-scale), for link prediction method using the best similarity measure, when attempting to predict $10^{-2}|E|$ to $0.1|E|$ unobserved edges $E^{U}$, for each graph in the dataset. For each similarity measure outlined in Section 3.1, we attempt only the best hub limit $L_{H}$ parameter setting obtained in Section 4.3 (for the LHub approach), and then select the best among them, considering both the F1 score and runtime. Note that the numerical suffix added to the acronym of each link prediction method, with the LHub approach, indicates the hub limit $L_{H}$ parameter setting.
@@
Figure 5. Overall phase split of the LHub approach, when predicting $0.1|E|$ links with similarity measures given in Section 3.1, while choosing suitable hub limit $L_{H}$ setting for each similarity measure as given in Section 4.3.
@@
Figure 6. Overall speedup of LHub approach for link prediction, and its phases (obtaining edges with top-k scores per thread, and merging scores from each thread into a common scoreboard), with $10^{-2}|E|$ unobserved edges, with increasing number of threads (in multiples of 2). Increasing the number of threads causes the work in merging phase to increase, thus leading to a poor speedup.
@@
As seen in Figure 4, the LHub approach is, on average, over 963×963$\times$963 × and 163×163$\times$163 × faster than the IHub approach with $10^{-2}|E|$ and $0.1|E|$ unobserved edges, respectively. It achieves this speedup while predicting links with an average F1 score that is $61\%$ higher and $16\%$ lower, respectively — meaning similar F1 scores without being too low or high. Notably, on the sk-2005 graph with $0.1|E|$ edges removed, LHub achieves a link prediction rate of $38.1M$ edges/s.
@@
Furthermore, we observe that link prediction with the RA metric excels, in terms of both F1 score and runtime, on web graphs and social networks when using the IHub approach. Meanwhile, link prediction with the JC similarity metric outperforms others on road networks and protein k-mer graphs. With the LHub approach and $10^{-2}|E|$ unobserved edges, link prediction with the CN metric (hub limit $L_{H}$ of $32$) excels on web graphs and social networks. For road networks, link prediction with the LHN metric (hub limit $L_{H}$ of $4$) performs the best, and for protein k-mer graphs, link prediction with the JC metric (hub limit $L_{H}$ of $256$) is optimal. However, with $0.1|E|$ unobserved edges, link prediction with the CN metric (hub limit $L_{H}$ of $32$) proves to be the best across all graphs. Figures 8, 9, 10, 11 show the runtimes and F1 scores for link prediction with all similarity measures. Notably, in Figures 10 and 11, LHub approach using the AA and RA metrics performs similarly to the CN metric, but with longer runtimes.
@@
Next, we note that the IHub approach achieves an average F1 score of $2.0\times 10^{-2}$ and $9.2\times 10^{-2}$ when predicting $10^{-2}|E|$ and $0.1|E|$ edges, respectively. In comparison, the LHub approach achieves higher F1 scores, averaging $3.2\times 10^{-2}$ and $9.8\times 10^{-2}$, respectively. Additionally, we observe that the F1 score tends to be higher on web graphs and social networks but significantly lower on road networks and protein k-mer graphs. This discrepancy is likely due to the average degree of the graphs, as local/neighborhood-based link prediction methods rely on the neighborhood of vertices up to a distance of $2$. Graphs with lower average degree provide less information to such methods for predicting edges. While these F1 scores may seem low (compared to the highest possible F1 score of $1$), it’s important to consider that they are significantly higher than random predictions. Although machine learning-based approaches might achieve higher F1 scores, neighborhood-based similarity measures excel in computational efficiency (both in terms of runtime and space) and interpretability, as mentioned earlier.

### v3

@@ Abstract.
Link prediction can help rectify inaccuracies in various graph algorithms, stemming from unaccounted-for or overlooked links within networks. However, many existing works use a baseline approach, which incurs unnecessary computational costs due to its high time complexity. Further, many studies focus on smaller graphs, which can lead to misleading conclusions. Here, we study the prediction of links using neighborhood-based similarity measures on large graphs. In particular, we improve upon the baseline approach (IBase), and propose a heuristic approach that additionally disregards large hubs (DLH), based on the idea that high-degree nodes contribute little similarity among their neighbors. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, DLH is on average 1019×1019$\times$1019 × faster than IBase, especially on web graphs and social networks, while maintaining similar prediction accuracy. Notably, DLH achieves a link prediction rate of $38.1M$ edges/s and improves performance by 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1.1. Our Contributions
In this report, we study parallel algorithms for efficient link prediction in large graphs using neighborhood-based similarity measures. First, we improve upon the baseline approach (IBase). It efficiently finds common neighbors and handles large graphs by tracking top-$k$ edges per-thread and later merges them globally. Next, we propose a novel heuristic approach that additionally discards large hubs (DLH), based on the observation that high-degree nodes contribute poorly to similarity among their neighbors. [^1] https://github.com/puzzlef/neighborhood-link-prediction-openmp We then determine suitable hub limits, i.e, the degree above which a vertex is considered a large hub, for link prediction with each similarity metric.
On a machine with two 16-core Intel Xeon Gold 6226R processors, our results show that the DLH approach outperforms IBase by over 1622×1622$\times$1622 × and 415×415$\times$415 × on average with $10^{-2}|E|$ and $0.1|E|$ unobserved edges, respectively. This speedup is achieved while maintaining comparable F1 scores. Notably, DLH achieves a link prediction rate of $38.1M$ edges/s with $0.1|E|$ unobserved edges. We also identified suitable similarity metric for each type of graph. When predicting $0.1|E|$ edges with the DLH approach, we observed that $63\%$ of the runtime is spent on the scoring phase, and especially higher on social networks, with high average degree. With doubling of threads, DLH exhibits an average performance scaling of 1.6×1.6$\times$1.6 ×.
@@ 4. Approach
4.1. Our Improved Baseline Approach (IBase)
@@ 4.1.3. Design of Hashtable
(a) Our Improved Baseline approach (IBase)
(b) Disregard hubs with degree $>8$ (DLH)
(c) Disregard hubs with degree $>4$ (DLH)
Figure 1. Illustration of our link prediction approach which Disregards Large Hubs (DLH), i.e., $1^{st}$ order neighbors with high degree. Here we focus on the neighborhood of vertex $1$, but the approach applies to each vertex in the graph. In the figure, $1$ is outlined in black, its $1^{st}$ order neighbors in red, its $2^{nd}$ order neighbors in blue, and explored/traversed vertices are shown with a yellow fill. Edge directions indicate traversal, with some $2^{nd}$ order vertices omitted for simplicity (dotted edges). (a) Depicts the our Improved Baseline (IBase) approach, which considers all $2^{nd}$ order neighbors of vertex $1$. (b) Presents our DLH approach, which considers only $2^{nd}$ order neighbors linked to $1$ through a small hub (degree $\leq 8$). This pruning reduces runtime and enhances prediction quality. (c) Illustrates DLH approach, where vertices with degree $>4$ are considered large hubs.
@@ 4.1.7. Parallelizing the process
We call this approach IBase. It has a time complexity of $O(ND^{2})$, where $N$ is the number of vertices in the graph, and $D$ is the maximum degree of a vertex in the graph.
4.2. Our Heuristic Approach which Disregards Large Hubs (DLH)
@@ 4.2.1. Explanation of the approach
We call this approach DLH. It has the same time complexity as IBase, i.e., $O(ND^{2})$, but significantly outperforms it by runtime, while achieving the same or better prediction quality.
@@ 4.2.2. A simple example
Figure 1 demonstrates the DLH approach. Here, we consider the neighborhood of vertex $1$ in a graph, where $1$ is outlined in black, its first-order neighbors in red, its second-order neighbors in blue, and explored/traversed vertices are shown with a yellow fill. The arrows in the figure indicate the direction of DFS traversal, and for figure simplicity, three neighbors of vertex $2$ are omitted (shown with dotted edges).
Figure 1(a) depicts the standard approach, which considers all second-order neighbors of vertex $1$ for score computation, i.e. vertices $6$ to $12$. This is the process followed by the IBase approach.
Figure 1(b) depicts our DLH approach. It considers only second-order neighbors linked to vertex $1$ by a small hub, with a degree $\leq 8$, for score computation — i.e., vertices $9$ to $12$ which are linked to $1$ through vertices $3$, $4$, and $5$. It is based on the idea, as mentioned above, that low-degree vertices contribute more similarity among their neighbors, while vertices with high degree do not. Note that second-order neighbors of $1$ can still be considered if linked to $1$ by a small hub (e.g., vertex $9$ linked to $1$ through $3$).
Lastly, Figure 1(c) shows the DLH approach, considering only second-order neighbors linked to vertex $1$ by a small hub with a degree $\leq 4$ ($11$ and $12$ linked to $1$ through $4$ and $5$). First order neighbors of $1$ with a degree $>4$ are considered large hubs, and thus their second-order neighbors are disregarded.
@@ 4.2.3. Non-determinism in the result
It’s worth noting that the results obtained by our parallel algorithms (IBase and DLH) are non-deterministic. This is because multiple edges can have matching scores, and the order of edges with the same score may be random due to parallelism.
@@ 4.3. Selecting Suitable Hub Limit for each Similarity Metric
We now select a suitable hub limit $L_{H}$, i.e., a degree above which a vertex is considered a large hub, for each similarity metric presented in Section 3.1. For this, we adjust the hub limit $L_{H}$ from $2$ to $1024$ (in multiples of $2$), for each similarity metric, and adjust the number of unobserved edges $E^{U}$ from $10^{-4}|E|$ to $0.1|E|$ on a number of graphs. We then apply each similarity metric based link prediction to predict $N_{P}=|E^{U}|$ edges with the highest similarity scores. We also test with a hub limit $L_{H}$ of $\infty$, which is the IBase approach.
Figure 3(a) shows the mean runtime taken for each similarity metric, with each hub limit $L_{H}$, and with the number of unobserved edges $E^{U}$ ranging from $10^{-4}|E|$ to $0.1|E|$, while Figure 3(b) shows the mean F1 score of the predicted edges, with the IBase and DLH approaches. Results indicate that selecting a lower hub limit $L_{H}$ decreases the overall runtime of the DLH approach. In terms of both F1 score and relative runtime, we observe that a hub limit $L_{H}$ of $4$ is suitable for HP and LHN metrics, a hub limit $L_{H}$ of $32$ is suitable for CN and AA metrics, and a hub limit $L_{H}$ of $256$ is suitable for the remaining metrics (i.e., JC, SI, SC, HD, and RA). These hub limits are highlighted with thick lines in the figures. Further, as Figure 7 shows, hub limits $L_{H}$ of $4$, $32$, and $256$ offer a mean speedup of 67×67$\times$67 ×, 32×32$\times$32 ×, and 13×13$\times$13 × when the number of unobserved edges $E^{U}$ is $0.1|E|$. Indeed, disregarding large hubs, with the DLH approach, offers a significant improvement in performance with little to no degradation in the accuracy of link prediction.
4.4. Our implementation of Disregard Large Hubs (DLH) approach
We now explain the implementation of our DLH approach for parallel neighborhood-based link prediction, which disregards large hubs. The pseudocode for this is given in Algorithm 1. Here, the predictLinks function accepts the input graph $G$, the maximum number of edges to predict $N_{P}$, a threshold score $S_{th}$, and outputs the list of predicted edges $P$. The algorithm operates in two phases: the scoring phase and the merging phase.
@@ 4.4. Our implementation of Disregard Large Hubs (DLH) approach
(a) Relative runtime (logarithmic scale) of each link prediction method
(b) F1 score of predicted links (logarithmic scale) of each link prediction method
Figure 3. Impact of adjusting the hub limit $L_{H}$ from $2$ to $1024$ (in multiples of $2$), and to $\infty$ (IBase approach), on the runtime (log-scale), and F1 score of predicted links (log scale), of each neighbor-based link prediction method, with the number of unobserved edges $E^{U}$ ranging from $10^{-4}|E|$ to $0.1|E|$. The full form of each link prediction method is given in Section 3.1.
Algorithm 1 Our Parallel Disregard Large Hubs (DLH) approach.
@@ 5.1.3. Dataset
Table 1. List of $13$ graphs obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019), with the directed graphs being marked with an asterisk. Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges and removing self-loops), and $D_{avg}$ is the average degree.
@@ 5.2. Comparative Performance Evaluation
In this section, we compare the performance of the DLH approach, which disregards large hubs, with the IBase approach (which does not). We conduct this comparison for observed graphs $E^{O}$ based on each graph in the dataset, with $10^{-2}|E|$ to $0.1|E|$ unobserved edges. This involves removing $10^{-2}|E|$ to $0.1|E|$ edges, with endpoints chosen uniformly at random (as explained in Section 5.1.4). We predict the same number of edges with both approaches, i.e., $N_{P}=10^{-2}|E|$ or $0.1|E|$. For each observed graph, we predict links using the IBase approach and the DLH approach with a suitable hub limit $L_{H}$ identified in Section 4.3. In Figures 4(a), 4(b), and 4(c), we plot the runtimes, speedups, and F1 scores, respectively, for only the best approach for each graph (considering both F1 score and runtime). In the figures, the labels indicate the abbreviations of the similarity metric used, followed by the value of the hub limit $L_{H}$ parameter setting (e.g., the label $CN32$ stands for the Common Neighbors (CN) similarity metric, with a hub limit $L_{H}$ of $32$, where first order neighbors with a degree greater than $32$ are avoided). It’s worth noting that a hub limit $L_{H}$ of $\infty$ essentially represents the IBase approach. Note that IBase approach crashed on sk-2005 graph with $0.1|E|$ unobserved edges due to out of memory issue, and thus these plots are not shown.
@@
(a) Runtime in seconds (logarithmic scale) for link prediction using the best similarity measure, with IBase and DLH approaches
@@
(b) Speedup (logarithmic scale) for link prediction with the best similarity measure of DLH approach, compared to IBase approach
@@
(c) F1 score of predicted links (logarithmic scale), for link prediction using the best similarity measure, with IBase and DLH approaches
@@
Figure 4. Runtime in seconds (log-scale), speedup (log-scale), and F1 score of predicted links (log-scale), for link prediction with the Improved Baseline (IBase) approach and our approach of Disregarding Large Hubs (DLH), using the best similarity measure, when attempting to predict $10^{-2}|E|$ to $0.1|E|$ unobserved edges $E^{U}$, for each graph. For each similarity measure outlined in Section 3.1, we attempt only the best hub limit $L_{H}$ parameter setting obtained in Section 4.3 (for the DLH approach), and then select the best among them, considering both the F1 score and runtime. Note that the numerical suffix added to the acronym of each link prediction method, with the DLH approach, indicates the hub limit $L_{H}$ parameter setting.
@@
Figure 5. Overall phase split of the DLH approach, when predicting $0.1|E|$ links with similarity measures given in Section 3.1, while choosing suitable hub limit $L_{H}$ setting for each similarity measure as given in Section 4.3.
@@
Figure 6. Overall speedup of our approach of Disregarding Large Hubs (DLH) for link prediction, and its phases (obtaining edges with top-k scores per thread, and merging scores from each thread into a common scoreboard), with $10^{-2}|E|$ unobserved edges, with increasing number of threads (in multiples of 2). Increasing the number of threads causes work in merging phase to increase, thus leading to a poor speedup.
@@
As seen in Figure 4, the DLH approach is, on average, over 1622×1622$\times$1622 × and 415×415$\times$415 × faster than the IBase approach with $10^{-2}|E|$ and $0.1|E|$ unobserved edges, respectively. It achieves this speedup while predicting links with an average F1 score that is $80\%$ higher and $13\%$ lower, respectively — meaning similar F1 scores without being too low or high. Notably, on the sk-2005 graph with $0.1|E|$ edges removed, DLH achieves a link prediction rate of $38.1M$ edges/s.
@@
Furthermore, we observe that link prediction with the RA metric excels, in terms of both F1 score and runtime, on web graphs and social networks when using the IBase approach. Meanwhile, link prediction with the JC similarity metric outperforms others on road networks and protein k-mer graphs. With the DLH approach and $10^{-2}|E|$ unobserved edges, link prediction with the CN metric (hub limit $L_{H}$ of $32$) excels on web graphs and social networks. For road networks, link prediction with the LHN metric (hub limit $L_{H}$ of $4$) performs the best, and for protein k-mer graphs, link prediction with the JC metric (hub limit $L_{H}$ of $256$) is optimal. However, with $0.1|E|$ unobserved edges, link prediction with the CN metric (hub limit $L_{H}$ of $32$) proves to be the best across all graphs. Figures 8, 9, 10, 11 show the runtimes and F1 scores for link prediction with all similarity measures. Notably, in Figures 10 and 11, DLH approach using the AA and RA metrics performs similarly to the CN metric, but with longer runtimes.
@@
Next, we note that the IBase approach achieves an average F1 score of $1.8\times 10^{-2}$ and $1.1\times 10^{-1}$ when predicting $10^{-2}|E|$ and $0.1|E|$ edges, respectively. In comparison, the DLH approach achieves F1 scores, averaging $3.2\times 10^{-2}$ and $9.8\times 10^{-2}$, respectively. Additionally, we observe that the F1 score tends to be higher on web graphs and social networks but significantly lower on road networks and protein k-mer graphs. This discrepancy is likely due to the average degree of the graphs, as local/neighborhood-based link prediction methods rely on the neighborhood of vertices up to a distance of $2$. Graphs with lower average degree provide less information to such methods for predicting edges. While these F1 scores may seem low (compared to the highest possible F1 score of $1$), it’s important to consider that they are significantly higher than random predictions. Although machine learning-based approaches might achieve higher F1 scores, neighborhood-based similarity measures excel in computational efficiency (both in terms of runtime and space) and interpretability, as mentioned earlier.